### v2

# Skyrme-Hartree-Fock-Bogoliubov mass models on a 3D mesh: IV. Improved description of the isospin dependence of pairing

Abstract
Providing reliable data on the properties of atomic nuclei and infinite nuclear matter to astrophysical applications remains extremely challenging, especially when treating both properties coherently within the same framework. Methods based on energy density functionals (EDFs) enable manageable calculations of nuclear structure throughout the entire nuclear chart and of the properties of infinite nuclear matter across a wide range of densities and asymmetries. To address these challenges, we present BSkG4, the latest Brussels-Skyrme-on-a-Grid model. It is based on an EDF of the extended Skyrme type with terms that are both momentum and density-dependent, and refines the treatment of ${}^{1}S_{0}$ nucleon pairing gaps in asymmetric nuclear matter as inspired by more advanced many-body calculations. The newest model maintains the accuracy of earlier BSkGs for known atomic masses, radii and fission barriers with rms deviations of 0.633 MeV w.r.t. 2457 atomic masses, 0.0246 fm w.r.t. 810 charge radii, and 0.36 MeV w.r.t 45 primary fission barriers of actinides. It also improves some specific pairing-related properties, such as the ${}^{1}S_{0}$ pairing gaps in asymmetric nuclear matter, neutron separation energies, $Q_{\beta}$ values, and moments of inertia of finite nuclei. This improvement is particularly relevant for describing the $r$-process nucleosynthesis as well as various astrophysical phenomena related to the rotational evolution of neutron stars, their oscillations, and their cooling.
Keywords:
energy density functional pairing neutron star superfluidity
journal: Eur. Phys. J. A
1 Introduction
Nuclear physics predictions are essential to many astrophysics applications, including simulations of the rapid neutron capture process (or $r$-process) and the modeling of neutron stars (NS). However, accurately describing the properties of the thousands of atomic nuclei involved in the $r$-process or of nuclear matter remains an arduous task, further complicated by the wide range of densities and neutron-proton asymmetries found in astrophysical environments Arnould20 ; Oertel17 ; Burgio18 . The challenge lies in developing a model that can simultaneously describe the equation of state (EoS) of dense matter in NS Oertel17 and nuclear structure properties (such as binding energies, nuclear level densities, fission probabilities, and photon or beta strength functions) that are crucial for modeling reactions and decays Arnould20 .
Methods based on energy density functionals (EDFs) provide an effective tool for supplying nuclear data to astrophysics, offering a quantum description of nuclear systems in terms of neutrons and protons Bender03 . By connecting nucleonic densities to an effective nucleon-nucleon interaction, EDFs allow for tractable calculations of nuclear structure across the entire nuclear chart and of infinite nuclear matter (INM) properties for a wide range of densities. This makes EDFs a powerful framework for advancing global nuclear models relevant to astrophysics. Within this framework, we have been developing the Brussels-Skyrme-on-a-Grid (BSkG) Scamps21 ; Ryssens22 ; Grams23 series of global nuclear structure models. The BSkG family is based on EDF of extended Skyrme type with both momentum- and density-dependent $t_{4}$ and $t_{5}$ terms, and uses the Hartree-Fock-Bogoliubov (HFB) approximation to address the many-body problem. These models aim to provide a comprehensive and as microscopic as possible description of nuclear structure and INM, with a specific focus on astrophysical applications. The BSkG series exploits the powerful concept of symmetry breaking, allowing for complex nuclear shapes, such as triaxial and octupole deformations at ground-state Grams23 and along the fission path Ryssens23 . The description of deformed nuclear clusters is also of importance for the deep layers of the NS inner crust, where a competition between the strong and the electromagnetic interactions takes place, leading to exotic nuclear configurations Ravenhall83 ; Hashimoto84 . These so-called nuclear pasta phases may look like “gnocchi" (droplets), “spaghetti” (rods), “lasagna” (slabs), “bucatini” (tubes), and “Swiss cheese” (bubbles) and are predicted to exist in cold and hot stellar matter. Moreover, BSkG models allow for time-reversal breaking and can hence access spin and current densities; the latter enter so-called ‘time-odd terms’ in the EDF, which contribute to the total energy of odd and odd-odd nuclei and are crucial for describing spin- and spin–isospin polarized INM. Aside from their application to $r$-process nucleosynthesis Goriely2020 , the BSkG models have shown how relevant their symmetry-breaking approach is to interpret new experimental data in different regions of the nuclear chart: from mass measurements hukkanen2023 ; hukkanen2023a ; Hukkanen22 to charge radii extracted from laser spectroscopy maass_inprep ; warbinek2024 and even the interpretation of newly discovered isomers stryjczykDiscoveryNewLonglived2024 .
In this work we concentrate on improving the ${}^{1}S_{0}$ pairing channel of our models, which plays an essential role in both finite nuclei and dense nuclear matter; pairing contributes to binding energy and stability of the former and leads to the emergence of superfluidity and superconductivity in the dense matter of NSs Chamel17 . Pairing arises due to an attractive interaction between fermions, which allows them to form pairs and condense into a superfluid state below a critical temperature Bardeen57 . Odd-even effects associated with the pairing interaction directly influence the neutron separation energies of finite nuclei, a quantity that defines the nuclear path during the $r$-process neutron irradiation, as well as their $Q_{\beta}$ value that determines the $\beta$-decay half-lives, hence the timescales against the production of heavy species during the $r$-process Arnould07 . Similarly, pairing effects are known to influence the prediction of fission probabilities that can affect the recycling and final composition resulting from the $r$-process nucleosynthesis Goriely15 . For these reasons, a careful description of the pairing interaction is needed for a reliable prediction of absolute masses as well as mass differences and fission properties influencing reaction and decay rates, hence nucleosynthesis simulations.
Another interesting physical system to the study of the pairing phenomena is the inner crust of a NS, where nuclear clusters coexist with superfluid neutrons and possibly superconducting protons in the deepest layers. Because the coherence lengths are typically large Okihashi2021 , a consistent and reliable treatment of the pairing phenomenon in both clusters and INM is required for an accurate modeling of superfluidity in the inner crust of NS Chamel2010super . Superfluidity plays a major role in various astrophysical phenomena, such as pulsar frequency glitches antonopoulou2022 , NS cooling potekhin2020 and oscillations Andersson21 . A new toolkit has been recently released Pecak24 to simulate the local superfluid dynamics in NS, allowing for instance to analyze the motion of an impurity through the neutron superfluid in the NS inner crust, and extract the effective mass, determine the critical velocities, study the formation of topological defects and dissipation processes. The tool is based on time-dependent HFB calculations with extended Skyrme functionals such as the BSkG series and will benefit from a more refined treatment of INM pairing gaps at arbitrary asymmetries in the EDF.
The last model of our series, BSkG3 Grams23 , describes nucleon pairing at low densities based on a specific prescription built to match the ${}^{1}S_{0}$ pairing gaps in INM as determined by realistic Extended Brueckner-Hartree-Fock (EBHF) calculations Cao06b , rather than relying on the purely phenomenological ansatz used in BSkG1-2 Scamps21 ; Ryssens22 . However, this match to EBHF calculations is limited to pure neutron matter (NM) and symmetric nuclear matter (SM), requiring an interpolation scheme for arbitrary asymmetries. All Brussels models since BSk17 Goriely09 (with the exception of BSk27 goriely2013 ) adopted a simple prescription that later was found to lead to negative proton pairing gaps at high asymmetries. We modified the interpolation prescription for BSkG3 Grams23 to avoid this issue. However, we found that the improvement on BSkG3 was not enough to reproduce the behaviour of pairing gaps at arbitrary asymmetries as given by Bardeen-Cooper-Schrieffer (BCS) approximation Zhang_Cao_etal_10 . To address this problem we propose a new entry for the BSkG series, which improves the description of pairing gaps in highly asymmetric environments. This is achieved with BSkG4 by modifying the pairing term of the EDF with an improved interpolation ansatz for the pairing gaps, while keeping the other model ingredients similar to those used for BSkG3, as detailed in Sec. 2. BSkG4 description of finite nuclei is discussed in Sec. 3 and its predictions for INM properties and astrophysics applications in Sec. 4. Finally, we reserve the last section to our conclusions.
2 Model construction and ${}^{1}S_{0}$ pairing gaps
We describe an atomic nucleus using a Bogoliubov type mean-field state, which we obtain as the solution to the self-consistent HFB equations based on a Skyrme-type EDF. BSkG4 was constructed along the same lines as BSkG3: it is based on an extended form of the Skyrme functional whose parameters resulted from a large-scale fitting protocol featuring thousands of known nuclear masses, hundreds of charge radii, reference values of the fission barriers of actinide nuclei and INM properties Grams23 . We relied on the efficient and robust MOCCa code; its three-dimensional coordinate-space representation imposes little symmetry restrictions on the nuclear configurations while providing excellent numerical precision, even for the elongated shapes encountered along fission barriers Ryssens15 ; RyssensThesis . We refer the interested reader for more information to Refs. Scamps21 ; Ryssens22 ; Grams23 .
The novelty of BSkG4 concerns our treatment of the pairing channel away from the limits of NM and SM. The pairing terms in our functional are inspired by those of a zero-range pairing interaction acting between like particles but with an effective density-dependent strength $g_{q}(\rho_{n},\rho_{p})$ ($q=p,n$):
$g_{q}(\rho_{n},\rho_{p})$ $=V_{q}(\rho_{n},\rho_{p})\left[1+\kappa_{q}(\nabla\rho)^{2}\right]\,,$ (1)
where the $\rho_{n},\rho_{p}$ are the nucleonic densities and $\rho=\rho_{n}+\rho_{p}$ is the total density. Using the procedure of Ref. goriely+2016 , we determine $V_{q}(\rho_{n},\rho_{p})$ by requiring that the pairing gaps corresponding to Eq. (1) locally match those of advanced many-body calculations in homogeneous INM. The ansatz for the function $g_{q}$ and the regularization of high-momentum divergences are the same as for BSkG3 Grams23 . In particular, this prescription allows for both the bulk and gradient pairing terms to be properly renormalized with the energy cutoff $E_{\rm cut}$. This is crucial for time-dependent HFB calculations pecak_2021 for which very high cutoff values $E_{\rm cut}\sim 100$ MeV have to be considered to ensure the conservation of energy during the evolution (see, e.g., Ref. magierski2019nuclear ).
We rely on the predictions for the ${}^{1}S_{0}$ nucleon pairing gaps in INM of Ref. Cao06 based on EBHF with polarization and self-energy corrections [^1] Recently we have implemented a new feature in the MOCCa code - an option for the strictly periodic boundary conditions. This allows us to verify that the code reproduces the reference gaps of Ref. Cao06 in NM. The detailed methods and results will be reported in a forthcoming paper.. In practice, we adopt parameterizations of those predictions; these are described in detail in B. Since the EBHF calculations were restricted to NM and SM, we are forced to seek an interpolation formula that can reasonably well describe pairing gaps for arbitrary isospin asymmetries $\delta=(\rho_{n}-\rho_{p})/\rho$. A reasonable initial step on this road is assuming isospin symmetry, i.e. that $\Delta_{n}(\rho_{n},\rho_{p})=\Delta_{p}(\rho_{p},\rho_{n})$. The first interpolation was proposed in Ref. Goriely09 :
$\Delta_{q}(\rho_{n},\rho_{p})=(1-|\delta|)\Delta_{\mathrm{SM}}(\rho)\pm\delta\frac{\rho_{q}}{\rho}\Delta_{\mathrm{NM}}(\rho_{q})\,,$ (2)
where the upper (lower) sign is to be taken for neutrons (protons). This interpolation was adopted in all of the Bogoliubov-Skyrme (BSk) parametrizations since BSk17 Goriely09 , except for BSk27 goriely2013 . However, Eq. (2) has the shortcoming that it produces negative pairing gaps in specific density ranges when $\delta$ was close (but not equal to) -1. Although one can manually set the gap to zero in such cases, we avoided this spurious feature for BSkG3 Grams23 by using instead:
$\Delta_{q}(\rho_{n},\rho_{p})=(1-|\delta|)\Delta_{\mathrm{SM}}(\rho)+|\delta|\Delta_{\mathrm{NM}}(\rho_{q})\,.$ (3)
However, Eq.(3) remains a simple ad-hoc proposal.
Zhang et al. Zhang_Cao_etal_10 investigated the isospin dependence of the ${}^{1}S_{0}$ pairing gaps in asymmetric INM from BCS calculations using the realistic Argonne AV18 potential. Although they did not include polarization and self-energy corrections, their results can still provide some guidance. As shown in Fig. 1, neither Eq. (2) (dotted lines in the upper panel) nor Eq. (3) (dashed lines in the middle panel) can reproduce the results of Ref. Zhang_Cao_etal_10 (this flaw of Eq. (2) was already pointed out in Ref. Zhang_Cao_etal_10 ). For this comparison, we use analytic representations for $\Delta_{\mathrm{NM}}$ and $\Delta_{\mathrm{SM}}$ fitted to the results from Ref. Zhang_Cao_etal_10 in NM and SM respectively.
Figure 1: ${}^{1}S_{0}$ neutron pairing gaps in asymmetric INM as a function of the isospin asymmetry parameter $\delta$ (the proton pairing gaps are the same if one changes $\delta$ to $-\delta$). Points represent the results of Ref. Zhang_Cao_etal_10 at different total densities $\rho$ (in fm-3). The dotted lines on the top panel correspond to the interpolation from Eq. (2). The middle panel shows the same quantity in dashed lines for the interpolation scheme of Eq. (3). The bottom panel displays in solid lines the new interpolation of Eq. (6). All three panels interpolate the symmetric and neutron matter gaps from Ref. Zhang_Cao_etal_10 .
This disagreement motivated us to search for a better interpolation. As was argued in Ref. Zhang_Cao_etal_10 , the isospin dependence of the pairing gaps mainly comes from the isospin splitting of the effective mass $m^{*}_{q}$. Therefore we adopt the ansatz
$\Delta_{q}(\rho_{n},\rho_{p})=\Delta_{\mathrm{NM}}(\rho_{q})\cdot\exp\left(\frac{C_{1}(\rho)}{m^{*}_{q}(\rho,\delta)}+C_{2}(\rho)\right).$ (4)
Then we can determine the functions $C_{1}(\rho)$ and $C_{2}(\rho)$ by requiring $\Delta_{n}(\rho,0)=\Delta_{p}(0,\rho)=\Delta_{\mathrm{NM}}(\rho)$ and $\Delta_{q}(\rho/2,\rho/2)=\Delta_{\mathrm{SM}}(\rho)$. Furthermore, the effective mass for the extended Skyrme form we use can be expressed as
$\dfrac{m^{*}_{q}(\rho,\delta)}{m_{q}}=\left[1+a(\rho)\pm b(\rho)\delta\right]^{-1}\,,$ (5)
with upper (lower) sign for $q=n(p)$ Chamel09 ; $a(\rho),\,b(\rho)$ are functions of the total density only. After some algebra, we find
$\Delta_{q}(\rho_{n},\rho_{p})$ $=\Delta_{\rm NM}(\rho_{q})\left[\frac{\Delta_{\rm SM}(\rho)}{\Delta_{\rm NM}(\rho/2)}\right]^{\left(1\pm\delta\right)}\,,$ (6)
with the lower (upper) sign for neutrons (protons). We show in the bottom panel of Fig. 1 that Eq. (6) is much better at capturing the isospin dependence of the predictions of Ref. Zhang_Cao_etal_10 .
We now apply the new formula to the EBHF predictions of Cao et al. Cao06 ; Fig. 2 compares the results with those produced by the interpolation used for BSkG3 (Eq. 3). Our new approach generally predicts smaller values of $\Delta_{n}$ in neutron-rich matter. We will see in Sec. 3 that this results in somewhat smaller pairing gaps in heavy nuclei compared to our previous interpolation. The largest differences, however, arise in proton-rich matter: $\Delta_{n}$ is enhanced at densities $\rho\lesssim 0.04$ fm-3 and suppressed at higher densities. As will be discussed in Sec. 4, the latter property has important consequences for proton superconductivity in the NS interior. Finally, the new interpolation removes the kink at $\delta=0$ and renders the derivation of thermodynamic quantities easier.
Figure 2: ${}^{1}S_{0}$ neutron pairing gaps in asymmetric INM for the BSkG3 (dashed lines) and BSkG4 (solid lines) parametrizations. Colors correspond to different total densities $\rho$ (in fm-3).
3 Properties of atomic nuclei
Table 1: The BSkG4 parameter set: seventeen parameters determining the self-consistent mean-field energy $E_{\rm HFB}$, three to the pairing functional, and nine determining the correction energy $E_{\rm corr}$ (see Ref. Grams23 for details on $E_{\rm HFB}$ and $E_{\rm corr}$). For comparison, we include the values of the BSkG3 parameter set Grams23 . Note that instead of parameter $x_{2}$ we list the values of the product $x_{2}\,t_{2}$.
[…]
$t_{0}$ [MeV fm3] $-2325.35$ $-2325.45$
$t_{1}$ [MeV fm5] $749.82$ $731.84$
$t_{2}$ [MeV fm5] $0.01$ $0.01$
$t_{3}$ [MeV fm3+3α] $14083.45$ $14092.79$
$t_{4}$ [MeV fm5+3β] $-498.01$ $-476.32$
$t_{5}$ [MeV fm5+3γ] $266.52$ $271.19$
$x_{0}$ $0.558834$ $0.549106$
$x_{1}$ $2.940880$ $2.97317$
$x_{2}t_{2}$ [MeV fm5] $-432.256954$ $-431.435904$
$x_{3}$ $0.628699$ $0.618431$
$x_{4}$ $5.657990$ $5.87636$
$x_{5}$ $0.396933$ $0.353345$
$W_{0}$ [MeV fm5] $119.735$ $122.206$
$W_{0}^{\prime}$ [MeV fm5] $78.988$ $79.840$
$\alpha$ $1/5$ $1/5$
$\beta$ $1/12$ $1/12$
$\gamma$ $1/4$ $1/4$
$\kappa_{n}$ [fm8] $123.20$ $123.20$
$\kappa_{p}$ [fm8] $129.07$ $129.07$
$E_{\rm cut}$ [MeV] $7.961$ $7.919$
$b$ $0.810$ $0.905$
$c$ $7.756$ $6.764$
$d$ $0.289$ $0.234$
$l$ $5.499$ $1.787$
$\beta_{\rm vib}$ $0.827$ $0.866$
$V_{W}$ [MeV] $-1.716$ $-1.411$
$\lambda$ $437.20$ $560.00$
$V_{W}^{\prime}$ [MeV] $0.502$ $0.531$
$A_{0}$ $37.801$ $38.174$
Table 2: Root-mean-square ($\sigma$) and mean ($\bar{\epsilon}$) deviations between experiment and predictions for the BSkG3 and BSkG4 models. The first block refers to the nuclear ground-state properties and the second one to fission properties. More specifically, these values were calculated with respect to 2457 known masses ($M$AME2020 of nuclei with $Z$, $N\geq 8$, the subset of 299 known masses $M_{nr}$ of neutron-rich nuclei with $S_{n}\leq 5$ MeV, the 1784 (2109) five-point neutron (proton) gaps $\Delta^{(5)}_{n}$ $(\Delta^{(5)}_{p})$ , 2309 neutron separation energies ($S_{n}$), 2173 $\beta$-decay energies ($Q_{\beta}$), 810 measured charge radii ($R_{c}$) Angeli13 , 45 empirical values for primary ($E_{\rm I}$) and secondary ($E_{\rm II}$) fission barrier heights Capote09 and 28 fission isomer excitation energies ($E_{\rm iso}$) of actinide nuclei Samyn04 . The first line gives the model error Moller88 on all the measured masses.
[…]
$\sigma_{\rm mod}(M)$ [MeV] $0.627$ $0.629$
$\sigma(M)$ [MeV] $0.631$ $0.633$
$\bar{\epsilon}(M)$ [MeV] $+0.080$ $+0.104$
$\sigma(M_{\rm nr})$ [MeV] $0.660$ $0.672$
$\bar{\epsilon}(M_{\rm nr})$ [MeV] $-0.011$ $+0.037$
$\sigma(\Delta_{n}^{5})$ [MeV] $0.300$ $0.290$
$\bar{\epsilon}(\Delta_{n}^{5})$ [MeV] $-0.074$ $-0.041$
$\sigma(\Delta_{p}^{5})$ [MeV] $0.428$ $0.427$
$\bar{\epsilon}(\Delta_{p}^{5})$ [MeV] $+0.005$ $+0.005$
$\sigma(S_{n})$ [MeV] $0.442$ $0.402$
$\bar{\epsilon}(S_{n})$ [MeV] $+0.009$ $+0.010$
$\sigma(Q_{\beta})$ [MeV] $0.534$ $0.493$
$\bar{\epsilon}(Q_{\beta})$ [MeV] $+0.021$ $+0.026$
$\sigma(R_{c})$ [fm] $0.0237$ $0.0246$
$\bar{\epsilon}(R_{c})$ [fm] $+0.0006$ $+0.0006$
$\sigma(E_{\rm I})$ [MeV] $0.33$ $0.36$
$\bar{\epsilon}(E_{\rm I})$ [MeV] $+0.06$ $-0.02$
$\sigma(E_{\rm II})$ [MeV] $0.51$ $0.53$
$\bar{\epsilon}(E_{\rm II})$ [MeV] $+0.01$ $-0.02$
$\sigma(E_{\rm iso})$ [MeV] $0.36$ $0.33$
$\bar{\epsilon}(E_{\rm iso})$ [MeV] $-0.05$ $+0.05$
BSkG4 shares all model ingredients with BSkG3, with the exception of the interpolation of pairing gaps in INM. Both models have therefore the same number of 29 parameters. The BSkG4 parameters were refitted using the same fitting protocol as for BSkG3 and starting with BSkG3 values, resulting in similar or identical values, as shown in Table 1. Hence, differences between both models arise mainly from the novel approach to the pairing channel. We refer the reader to Refs. Scamps21 ; Ryssens22 ; Grams23 for more information on the role of all parameters and the fitting protocol.
3.1 Ground state properties
Table 2 showcases BSkG4 ability to globally describe the properties of atomic nuclei. The tables first block concerns ground state properties. We list the mean (ϵ¯=¯italic-$\bar{\epsilon}=$ experiment - model) and rms ($\sigma$) deviations of the model compared to AME20 data AME2020 for nuclei with $N,Z\geq 8$: their nuclear masses ($M$), separation energies ($S_{n}$), five-point neutron and proton gaps $(\Delta^{(5)}_{n/p})$ and $\beta$-decay energies $(Q_{\beta})$; we also show the mean and rms deviations of the charge radii we obtain w.r.t. 810 measured charge radii of Ref. Angeli13  [^2] We exclude the values of Re, Po, Rn, Fr, Ra, Am and Cm isotopes listed in Ref. Angeli13 : these were determined from systematics rather than measurements.. In the second block of Table 2, we treat the fission properties of actinide nuclei: the mean and rms deviations with respect to all 45 reference values in the RIPL-3 database Capote09 for the primary ($E_{\rm I}$) and secondary $(E_{\rm II})$ barriers of $Z\geq 90$ nuclei are listed, as are the mean and rms deviations of the excitation energies of 28 known fission isomers among the same set of nuclei.
Generally speaking, BSkG4 achieves the same level of global agreement with data as BSkG3: the rms deviations for the masses for the two models are nearly identical, both globally as well as when restricted to the subset of 299 known masses of neutron-rich nuclei with $S_{n}\leq 5$ MeV. This is reflected too in Fig. 3, which shows the difference between the calculated and experimental masses as a function of neutron number (top panel) and proton number (bottom panel). This figure is comparable to its equivalent for BSkG3 (Fig. 2 in Ref. Grams23 ) and in particular exhibits the same patterns w.r.t. the location of the largest deviations. The difference between BSkG4 and BSkG3 atomic masses is shown in Fig. 4 for all the 6946 nuclei with $8\leq Z\leq 118$ lying between the BSkG4 proton and neutron drip lines. Both models produce very similar masses, with a difference typically below 1 MeV. Some differences appear for exotic heavy nuclei, where BSkG4 obtains binding energies predominantly larger than BSkG3.
Nevertheless, BSkG4 performs better for mass differences: $\sigma(S_{n})$ and $\sigma(Q_{\beta})$ are reduced by 9% and 8%, respectively. Although these improvements appear modest, they establish BSkG4 as the most performant of all BSk and BSkG models for these observables; we are not aware of any other set of microscopic calculations that can match this level of accuracy for all $S_{n}$ and $Q_{\beta}$. On the other hand, BSkG4 performs slightly worse than BSkG3 for the rms deviations of charge radii and the primary fission barriers of actinide nuclei; even if $\sigma(R_{c})$ and $\sigma(E_{I})$ increased by 4% and 9% respectively. The mean deviation of the $R_{c}$ however remains the same level as BSkG3, and $\bar{\epsilon}(E_{I})$ is three times lower than for BSkG3. In summary, BSkG4 retains a global description of both properties that is highly competitive.
Figure 3: Differences between experimental AME2020 and BSkG4 masses as a function of the neutron number (top) and proton number (bottom).
Figure 4: Differences between BSkG4 and BSkG3 Grams23 masses as a function of the neutron number (top) and proton number (bottom) for all the 6946 nuclei with $8\leq Z\leq 118$ lying between the BSkG4 proton and neutron drip lines.
The rms deviations of the five-point pairing gaps in Table 2 do not immediately point to an improvement in our description of nuclear pairing properties: $\sigma(\Delta_{n}^{(5)})$ was improved by only 3% while $\sigma(\Delta_{p}^{(5)})$ is essentially unchanged. The mean deviation of the neutron pairing gaps indicates a strong improvement though: $\bar{\epsilon}(\Delta^{(5)}_{n})$ was nearly cut in half. As foreshadowed in Sec. 2, what produces these changes is the modification to our effective neutron pairing strength in heavy nuclei: the new interpolation formula produces smaller neutron pairing gaps for modest positive asymmetry $\delta$ across the range of densities most relevant to pairing in finite nuclei. The improvement becomes more evident when one restricts the comparison to nuclei with $A\geq 150$; for that data set, $\epsilon(\Delta^{(5)}_{n})=-0.082$ MeV for BSkG3 and only $-0.029$ MeV for BSkG4. However, there is no global improvement for the proton five-point gaps.
We illustrate these features with two local examples: the top (bottom) panel of Fig. 5 shows the neutron (proton) five-point difference along the Pu isotopic ($N=140$ isotonic) chain. The $\Delta^{(5)}_{n}$ predicted by BSkG4 are significantly lower than those of BSkG3 along the entire $Z=94$ chain; for the small range of neutron numbers where experimental data is available, BSkG4 performs markedly better than its predecessor even if the experimental trend is not perfectly reproduced near $N\sim 150$. On the other hand, the match between the calculated $\Delta^{(5)}_{p}$ values and experiment along the $N=140$ chain is qualitative at best for both models, even if BSkG4 arguably grasps the average of the experimental values somewhat better. Changing the interpolation formula is clearly not sufficient to cure such gross mismatches between experimental and calculated trends; these likely indicate a more fundamental issue with the single-particle structure predicted for these isotopes by both models Bender00b .
Nuclear pairing impacts other observables besides the binding energy: one example is the rotational moment of inertia (MOI). This observable is sensitive to both pairing and deformation and is crucial to models like ours: it serves as an additional ingredient in the modelling of various collective effects that cannot be captured by simple mean-field calculations, from contributions to the binding energy Scamps21 to the rotational enhancement of nuclear level densities hilaire01 . Fig. 6 shows the calculated MOIs [^3] For simplicity, we show the calculated Belyaev MOIs multiplied by 1.32; we have confirmed in Ref. Ryssens22 that this old recipe of Ref. Libert99 is an acceptable substitute for the determination of Thouless-Valatin MOIs based on full-fledged cranking calculations. for both models as well as experimental data; the comparison concerns all even-even nuclei with known $2_{1}^{+}$ excitation energies below 150 keV. Although it performs reasonably well for medium-mass nuclei, BSkG3 systematically underestimates the MOI for nuclei in the actinide region as already noted in Ref. Grams23 , a deficiency shared with both of its predecessors Scamps21 ; Ryssens23 . By reducing the neutron pairing strength, the moments of inertia for heavy nuclei trend upwards such that BSkG4 now provides a globally acceptable description of rotational MOIs.
Finally, we comment on a striking feature of the new model: the prevalence of ground state octupole deformation. When including this degree of freedom, 245 nuclei gained more than 500 keV compared to a reflection-symmetric calculation as opposed to just 196 isotopes with BSkG3. The largest energy gain is consistently predicted for 286Bk, but the BSkG4 value (3.61 MeV) is much larger than the BSkG3 one (2.91 MeV). This is a direct result of the new pairing treatment: the smaller pairing strength ensures less competition for (weak) shell effects near the octupole magic numbers to create static ground state reflection asymmetry Butler96 ; Ryssens2019b ; Chen21 .
Figure 5: A comparison of five-point gaps $\Delta^{(5)}_{q}$ obtained with BSkG3 and BSkG4 to experimental values. Top panel: neutron values along the Pu isotopic chain. Bottom panel: proton values along the $N=140$ isotonic chain.
Figure 6: Calculated moments of inertia (MOI) as a function of neutron number for BSkG3 (red circles) and BSkG4 (blue squares). Experimental data Zeng94 ; Afanasjev00 ; Pearson91 ; Raman01 are shown by black crosses. See text for details.
3.2 Fission properties of BSkG4
In addition to the mean and rms deviations shown in Table 2, this section analyzes the differences between BSkG3 and BSkG4 in terms of fission properties. Fig. 7 displays the primary and secondary fission barriers (upper and middle panels, respectively) as well as the isomer excitation energy (lower panel) for the $Z\geq 90$ nuclei with known empirical barriers included in the RIPL-3 database Capote09 . The numerical details of the calculations performed with the new model are the same as those in the previous work on BSkG2 Ryssens23 .
In the upper panel of Fig. 7 we see that BSkG4 predicts slightly higher primary barriers for most nuclei, compared to BSkG3. This is directly attributable to the new pairing treatment: BSkG3 predicted lower primary barriers due to an overestimation of pairing effects. The differences become larger as the mass increases. for instance, Am and Cm chains present more deviations in comparison with Pa or U.
A similar pattern is found for the secondary barriers (middle panel of Fig.7). With the exception of the Th chain, BSkG4 generally produces higher values than BSkG3, with these differences becoming more pronounced as the neutron number increases. Although in some isotopic chains (e.g., Th or Cm) BSkG4 predictions deviate more from the reference values compared to BSkG3, the overall level of agreement remains comparable (see rms deviations in Table 2).
In the case of the isomer excitation energy, BSkG4 systematically predicts lower values. With the reduced pairing content, the shell effects become more pronounced compared to the previous interaction, leading to deeper wells along the fission paths.
Figure 7: Energies of the primary (upper panel) and secondary (middle panel) fission barriers and fission isomer (lower panel) for the 45 actinides in the RIPL-3 database obtained with BSkG3 (red dots) and BSkG4 (blue dots), along with the available reference values (black crosses) Capote09 . The grey stripes help to differentiate between the isotopic chains shown.
Overall, both models predict the same general topography for the actinide barriers. While the deviation between them is notable, it remains relatively small compared to the range of calculated values reported in the literature  Ryssens23 . Specifically, the mean and rms deviations between BSkG4 and BSkG3 for the 45 nuclei with $Z\geq 90$ are $\bar{\epsilon}=0.08$ MeV and $\sigma=0.16$ MeV, respectively, for the primary barrier, 0.03 MeV and 0.09 MeV for the secondary barrier and -0.07 MeV and 0.11 MeV for the isomer excitation energy. Hence, on average, BSkG4 barriers are higher than those of BSkG3 with deviations of the order of 0.1 MeV.
To rapidly test if such deviations between BSkG3 and BSkG4 may grow with isospin, we also analyzed the neutron-rich isotope 260U. Fig. 8 displays the projected 1-dimensional minimum-energy fission path as a function of $\beta_{20}$ for both BSkG3 and BSkG4 parameterizations. To account for triaxial effects, we computed the energy surface in the $(\beta_{20},\beta_{22})$ plane by imposing constraints on both $q_{20}$ and $q_{22}$. Once the surface was defined, we used the PyNEB code Flynn2022 to trace the minimum energy (one-dimensional) path from the ground state to the fission configuration.
Figure 8: Minimum energy fission paths of 260U as a function of $\beta_{20}$ obtained with BSkG3 and BSkG4.
In contrast to the trend observed in the U chain in Fig. 7, BSkG4 predicts slightly smaller primary and secondary barriers. In addition, BSkG4 shows a small initial hump around $\beta_{20}\approx 0.3$, whereas BSkG3 exhibits a flat region at this deformation. Despite these differences, both paths exhibit a similar overall topology with only minor deviations. The refit of the rotational correction parameters plays a significant role: increasing its value (see Table 1) mitigated the effect of higher rotational MOI induced by BSkG4 lower pairing.
Although we focused on a static description of fission, the predictions of BSkG4 do not differ significantly from those of BSkG3. The level of agreement with empirical values for actinides is similar, and for the specific case of the exotic neutron-rich 260U, the static fission paths show only small differences. The generalisation of this statement to all exotic neutron-rich nuclei and the impact of the larger MOI predicted by BSkG4 on the dynamical description of the fission path (where the action is minimized rather than the energy) will be explored in a future work.
4 Nuclear matter properties and astrophysics applications
4.1 Infinite nuclear matter properties
We present in Table 3 the properties of INM at saturation for BSkG3Grams23 and BSkG4. We show, from top to bottom, the Fermi momentum $k_{F}$, the saturation density $\rho_{\mathrm{sat}}$, the energy per particle of symmetric matter at saturation $a_{v}$, the symmetry energy coefficient $J$, the slope of the symmetry energy $L$, the isoscalar (isovector) effective mass $M_{s}^{*}/M$ ($M_{v}^{*}/M$), the compressibility modulus $K_{v}$, the isovector component of the compressibility coefficient $K_{\rm sym}$, the skewness parameter $K^{\prime}$ and the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ of symmetric INM. For BSkG4, we fix $J=31$ MeV, as previously done for BSkG3, since we can achieve a stiff symmetry energy thanks to the extended Skyrme form Grams23 . Note that both models present similar INM properties by construction, since we fitted BSkG4 starting from BSkG3. Both models present INM properties at saturation within the current uncertainties Margueron2018a .
Table 3: INM properties for the BSkG3 Grams23 and BSkG4 parameterisations. See Refs. Goriely16 ; Margueron02 ; Chamel10 for the various definitions.
[…]
$k_{F}$ [fm-1] $1.3270$ $1.3270$
$\rho_{\mathrm{sat}}$ [fm-3] $0.15784$ $0.15784$
$a_{v}$ [MeV] $-16.082$ $-16.079$
$J$ [MeV] $31.000$ $31.000$
$L$ [MeV] $55.061$ $55.674$
$M_{s}^{*}/M$ $0.85860$ $0.86046$
$M_{v}^{*}/M$ $0.72171$ $0.72748$
$K_{v}$ [MeV] $242.56$ $242.57$
$K_{\text{sym}}$ [MeV] $-21.248$ $-19.479$
$K^{\prime}$ [MeV] $304.17$ $303.94$
$G_{0}$ $0.20046$ $0.21786$
$G_{0}^{\prime}$ $0.97812$ $0.97888$
4.2 Neutron star properties
Table 4: NS properties obtained with BSkG3 and BSkG4 parameterizations. We show in the first row the radius of the canonical NS of $1.4M_{\odot}$; the following three lines give the mass, radius and energy density for the most compact NS. $\mathcal{E}_{\rm causal}$ denotes the energy density for which causality is violated. Next we show the threshold energy density and NS mass for the onset of the dUrca process and the energy density $\mathcal{E}_{\rm pQCD}$ up to which the model is consistent with constraints derived from pQCD (see text for details).
[…]
$R_{\rm 1.4}$ [km] $12.73$ $12.74$
$M_{\rm max}$ [$M_{\odot}$] $2.26$ $2.26$
$R_{\rm M_{\rm max}}$ [km] $11.10$ $11.09$
$\mathcal{E}_{\rm M_{\rm max}}$ [$10^{15}$ g cm-3] $2.273$ $2.282$
$\mathcal{E}_{\rm causal}$ [$10^{15}$ g cm-3] $2.381$ $2.384$
$\mathcal{E}_{\rm dUrca}$ [$10^{14}$ g cm-3] $7.136$ $7.083$
$M_{\rm dUrca}$ [$M_{\odot}$] $1.38$ $1.37$
$\mathcal{E}_{\rm pQCD}$ [$10^{15}$ g cm-3] $2.508$ $2.510$
[…]
The first three rows of Table 4 correspond to the radius of the canonical 1.4 $M_{\odot}$ NS, the maximum mass of NS and its respective radius, respectively. Next we show the energy density $\mathcal{E}_{\rm causal}$ of $\beta$-equilibrated matter for which the speed of sound exceeds the speed of light. We remark that $\mathcal{E}_{\rm causal}$ for BSkG4 is higher than the central energy density $\mathcal{E}_{\rm M_{\rm max}}$ of the most massive NS. Therefore the newest model (as BSkG3) does not violate causality.
To check if BSkG4 allows for the onset of the direct Urca (dUrca) process Lattimer91 we also show in Table 4 the density and the corresponding NS mass for which the proton fraction inside the star reaches the dUrca threshold. This process is required to interpret the observed thermal luminosity of some NS Burgio21 ; Marino2024 ; Ho_Pol_Deller_Becker_Burke-Spolaor_2024 . The BSkG4 and BSkG3 models allow for the dUrca process for NS with a mass above $\sim 1.4$ M.
To further test our models, we use results from perturbative quantum chromodynamics (pQCD) calculations: assuming only causality and thermodynamic consistency, pQCD predictions at extreme densities ( $\gtrsim$ 40 $\rho_{\rm sat}$) impose constraints on the EoS at lower densities Komoltsev22 . We have verified that the BSkG4 EoS is consistent with this constraint up to an energy density $\mathcal{E}_{\rm pQCD}=2.510\times 10^{15}$ g cm3, i.e. above the density of the most massive NS predicted by this model. BSkG3 also satisfies this constraint.
To illustrate the impact of BSkG4 model improvement on NS matter, we plot in Fig. 9 the proton pairing gap in $\beta$-equilibrated infinite $npe\mu$ matter. We compare BSkG4 (blue solid line), BSkG3 Grams23 (red dashed line) and BSk31 goriely+2016 (cyan short-dashed) with more advanced many-body calculations from: Baldo & Schulze 2007 (black dash-dash-dotted line) Baldo07 , Guo et al. 2019 (magenta crosses) Guo19 , and Lim & Holt 2021 (green dash-dotted line) obtained using the fitting formula for the n3lo450 plus second-order single-particle energy model from Lim21 with the proton fraction equal to the one of BSkG4. We show the density of crust-core transition with a vertical black line. In the core of the star, where protons are unbound, their pairing gap within BSkG4 tends closer to the realistic calculations given by Lim21 ; Baldo07 ; Guo19 . Even if the spread in the predictions of advanced many-body calculations indicates the uncertainty of the modelling of pairing in NM, the curves of BSkG3 Grams23 and BSk31 goriely+2016 are clearly too extreme.
[…]
In summary, the NS properties predicted by BSkG4 are almost identical to the ones obtained with BSkG3. The improvement of BSkG4 appears for properties sensitive to the pairing gaps, as shown in Fig. 9.
4.3 Simulation of $r$-process nucleosynthesis
[…]
Figure 10: Final mass fractions $X$ of the material ejected from a $1.375-1.375\leavevmode\nobreak\ M_{\odot}$ NS-NS binary system of Ref. Just23 , obtained with the new BSkG4 and former BSkG3 Grams23 masses. The solar system r-abundance distribution (open circles) from Goriely99 is shown for comparison and arbitrarily normalised. The lower panel shows the ratio between the predicted BSkG4 and BSkG3 mass fractions for mass fractions $X>10^{-8}$.
5 Conclusions
We presented a new entry in the BSkG series of large-scale nuclear structure models: instead of relying on an ad-hoc recipe, we developed a physically motivated approach to extend the EBHF results of Ref. Cao06 that guide our treatment of the EDFs pairing terms to arbitrary asymmetries. Compared to BSkG3, this modification leads to a modestly but systematically improved description of pairing in heavy nuclei and significantly different predictions for the pairing properties of neutron-rich matter in NS that qualitatively match more advanced many-body calculations. These improvements do not rely on the introduction of additional parameters nor do they come at the expense of our description of other quantities: the global rms deviations on known masses, charge radii and the fission properties of actinide nuclei are close to those of BSkG3, while the rms deviations of $S_{n}$ and $Q_{\rm\beta}$ have been lowered beyond those of any other microscopic model known to us. Furthermore, we performed $r$-process nucleosynthesis simulations with the new mass model. Our results show that the modifications with respect to the previous model have a non-negligible impact on the abundance distribution.
With its more physically motivated interpolation of ${}^{1}S_{0}$ pairing gaps between SM and NM, BSkG4 is better suited for modelling the thermal evolution of NS than any of our previous EDFs. BSkG4 will also allow for more realistic simulations of the neutron superfluid dynamics in NS, which remains poorly understood despite its importance for interpreting various astrophysical phenomena related to NS, such as glitches.
The improvement presented here have led to better description of pairing properties in both finite nuclei and INM. Nevertheless, the lack of ab initio data across the full range of asymmetries remains a key challenge. In particular, results that incorporate all relevant many-body effects, such as medium polarization and self-energy corrections, are sparse. We thus encourage further studies of the isospin dependence of the ${}^{1}S_{0}$ pairing gaps in INM. This would significantly advance both the developments of EDFs and their astrophysical applications.
Acknowledgements.
This work was supported by the Fonds de la Recherche Scientifique (F.R.S.-FNRS) and the Fonds Wetenschappelijk Onderzoek - Vlaanderen (FWO) under the EOS Projects nr O022818F and O000422F. This research benefited from computational resources made available on the Tier-1 supercomputer Lucia of the Fédération Wallonie-Bruxelles, infrastructure funded by the Walloon Region under the grant agreement nr 1117545. Further computational resources have been provided by the clusters Consortium des Équipements de Calcul Intensif (CÉCI), funded by F.R.S.-FNRS under Grant No. 2.5020.11 and by the Walloon Region. W.R. and S.G. gratefully acknowledge support by the F.R.S.-FNRS. This work also received funding from the Fonds de la Recherche Scientifique (Belgium) under Grant No. PDR T.004320 and IISN 4.4502.19.
Appendix A List of abbreviations
Abbreviation Meaning
BSkG Brussels-Skyrme-on-a-Grid
BCS Bardeen-Cooper-Schrieffer
dUrca direct Urca (process)
EBHF Extended Brueckner-Hartree-Fock
EDF Energy density functional
EoS Equation of state
GW Gravitational waves
HFB Hartree-Fock-Bogoliubov
INM Infinite (homogeneous) nuclear matter
MOI Moment of inertia
NM Pure (infinite) neutron matter
NS Neutron star
SM symmetric (infinite) nuclear matter
pQCD Perturbative quantum chromodynamics
rms root-mean-square
r-process rapid neutron-capture process
Appendix B Parameterization of the pairing gaps of Ref. Cao06
For ease of use in MOCCa, we parametrize the gaps $\Delta_{\mathrm{SM}}(\rho)$ and $\Delta_{\mathrm{NM}}(\rho_{q})$ as:
$\Delta_{\mathrm{NM}}(\rho_{q})=\theta(k_{m}-k_{Fq})\Delta_{0}\frac{k_{Fq}^{2}}{k_{Fq}^{2}+k_{1}^{2}}\,\frac{(k_{Fq}-k_{2})^{2}}{(k_{Fq}-k_{2})^{2}+k_{3}^{2}}$ (7)
and
$\Delta_{\mathrm{SM}}(\rho)=\theta(k_{m}-k_{F})\Delta_{0}\frac{k_{F}^{2}}{k_{F}^{2}+k_{1}^{2}}\,\frac{(k_{F}-k_{2})^{2}}{(k_{F}-k_{2})^{2}+k_{3}^{2}}\,.$ (8)
Here $k_{Fq}=(3\pi^{2}\rho_{q})^{1/3}$, $k_{F}=(3\pi^{2}\rho/2)^{1/3}$ and $\theta$ is the unit-step Heaviside function. The parameters $\Delta_{0},k_{1},k_{2},k_{3}$ and $k_{m}$ were determined by fitting the results of Ref. Cao06 ; the resulting values are listed in Table 5. These fits were performed for the construction of BSk30-31-32; although the parameter values were not provided at the time, the quality of the fit can be judged from Fig. 1 in Ref. goriely+2016 .
Table 5: Parameters for Eqs. (7)-(8) of the ${}^{1}S_{0}$ pairing gaps in SM and NM calculated in Ref. Cao06 .
SM NM
$\Delta_{0}$ [MeV] 11.5586 3.37968
$k_{1}$ [fm-1] 0.489932 0.556092
$k_{2}$ [fm-1] 1.31420 1.38236
$k_{3}$ [fm-1] 0.906146 0.327517
$k_{m}$ [fm-1] 1.31 1.38
Appendix C Explanation of the supplementary material
We provide as supplementary material the files
Mass_Table_BSkG4.dat and Fission_Table_BSkG4.dat. The former contains the calculated ground state properties of all nuclei with $8\leq Z\leq 118$ lying between the proton and neutron drip lines. The latter contains the fission barriers and isomer excitation energies as calculated for all 45 nuclei with $Z\geq 90$ that figure in the RIPL-3 database. The contents of both files follow the conventions of the supplementary files of Ref. Grams23 . For convenience, we repeat the contents of all columns of both files in Tables 6 and 7.
Column Quantity Units Explanation
1 Z -- Proton number
2 N -- Neutron number
3 $M_{\rm exp}$ MeV Experimental atomic mass excess
4 $M_{\rm th}$ MeV BSkG4 atomic mass excess
5 $\Delta M$ MeV $M_{\rm exp}-M_{\rm th}$
6 $E_{\rm tot}$ MeV Total energy
7 $\beta_{20}$- Quadrupole deformation
8 $\beta_{22}$-
9 $\beta_{2}$-
10 $\beta_{30}$- Octupole deformation
11 $\beta_{32}$-
12 $E_{\rm rot}$ MeV Rotational correction
13 $\langle\Delta\rangle_{n}$ MeV Average neutron gap
14 $\langle\Delta\rangle_{p}$ MeV Average proton gap
15 $r_{\rm BSkG4}$ fm Calculated rms charge radius
16 $r_{\rm exp}$ fm Experimental rms charge radius
17 $\Delta r$ fm $r_{\rm exp}-r_{\rm BSkG4}$
18 $\mathcal{I}^{B}$ 2superscriptPlanck-constant-over-$\hbar^{2}$ MeV-1 Calculated Belyaev MOI.
19 par(p) -- Parity of proton qp. excitation
20 par(n) -- Parity of neutron qp. excitation
Table 6: Contents of the Mass_Table_BSkG4.dat file.
Column Quantity Fission property Units Explanation
1 Z -- Proton number
2 N -- Neutron number
3 $E$ Inner barrier MeV Barrier height
4 $\beta_{20}$- Quadrupole deformation
5 $\beta_{22}$-
6 $E$ Outer barrier MeV Barrier height
7 $\beta_{20}$- Quadrupole deformation
8 $\beta_{22}$-
9 $\beta_{30}$- Octupole deformation
10 $E$ Isomer MeV Excitation energy
11 $\beta_{20}$- Quadrupole deformation
Table 7: Contents of the Fission_Table_BSkG4.dat file.